OpenStudy (anonymous):

A map has a scale of 1cm:20km. Two cities are 2.5 apart on the map. to the nearest tenth of a kilometer. What is the actual distance corresponding to the map distance? A.) 50km B.) 52.5km C.) 150km D.) 70km

pooja195 (pooja195):

1/20 = 2.5/x x = 50 km